Home‎ > ‎Crew 169 Team Peru IV‎ > ‎

0505-DailySummaryReport

DAILY SUMMARY REPORT
SOL: 4  5May2016
Person filling out Report: XO
Summary Title First sampling (Phase IIa)

Mission Status: Initial sampling.

Sol Activity Summary: Today I performed a first protocol for microbiological sampling (includes gloves, coat and bots sampling). In addition, Quality Control was performed according to CLSI M100-A2 guide.

Look Ahead Plan: Development of a new established sampling protocol. This includes a sampling in the exposed parts of coat, isolated in 3M PetriFlim media and incubation evaluation.

Anomalies in work: No problems.
Weather: First QC IIa.
Crew Physical Status: Adequate.
EVA: Initial (0).
Reports to be filed: Commander Report, XO Report, Engineering Report, HSO report and Daily Summary Report.
Equipment utilized: Oven at 37ºC, distilled water and volumetric equipment.